OpenStudy (anonymous):

Determine whether the integral converges or diverges. Find the value of the integral if it converges.

OpenStudy (anonymous):

What's the integral?

OpenStudy (anonymous):

\[\int_1^\infty x^{-1/3}~dx\] Write as a limit: \[\lim_{b\to\infty}\int_1^b x^{-1/3}~dx\] Then, like with the last one, apply the power rule and FTC: \[\lim_{b\to\infty}\left[\frac{x^{2/3}}{\frac{2}{3}}\right]_1^b\\ \frac{3}{2}\lim_{b\to\infty}\left[x^{2/3}\right]_1^b\\ \frac{3}{2}\left(\lim_{b\to\infty}b^{2/3}-1^{2/3}\right)\\ \]

OpenStudy (anonymous):

okay so it converges

OpenStudy (anonymous):

okay thanks

OpenStudy (anonymous):

No, it diverges. \[\lim_{b\to\infty}b^{2/3}=\infty\]

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

so that last one was diverges too since the lim didn't exist and it wasn't finite?

OpenStudy (anonymous):

Divergent.\[\int\limits_{1}^{\infty}x^{-\frac{1}{3}}dx>\int\limits_{1}^{\infty}x^{-1}dx\]Which is divergent.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

got you